OpenStudy (anonymous):

help me! 125 over 60 times five as a percent

OpenStudy (solomonzelman):

\[\frac{125}{60 \times 5}~~~or~~~\frac{125}{60}\times 5\]

OpenStudy (anonymous):

is it 41%? that's all I need to know

OpenStudy (solomonzelman):

is it the 1st thing or second from what I said?

OpenStudy (anonymous):

125/60 * 5/100

OpenStudy (anonymous):

first

OpenStudy (anonymous):

Is what the question mean is I guess

OpenStudy (solomonzelman):

\[Then~~~i t~~~is~~~~~~~~~~\frac{125}{300}\]

OpenStudy (solomonzelman):

in percent, 41.6666

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

Anytime!